# Efficient Testing Implies Structured Symmetry

### v1

@@ 3 Finding Structured Symmetry
In this section, we prove Theorems 1.3 and 1.2, which comprise our main equivalence between efficient testability and structured symmetry. As discussed in the introduction, this result strengthens the main theorem of [BY19] by capturing the relationship between the tester’s complexity and the structure of the property’s invariant sets. For convenience, we recall the statements of these two theorems below. The first, Theorem 1.2, is the harder of the two, showing that efficient testing implies structured symmetry.
@@
While the proof of Theorem 1.3 is quite simple, the proof of Theorem 1.2 is not, requiring two key lemmas. We call them the oracle simulation lemma and the tester simulation lemma, and we prove them in Sections 3.1 and 3.2, respectively. While both lemmas involve regular simulators, they differ in terms of which objects play the roles of the target function and distinguisher family. For the oracle simulation lemma, the target function is the one we wish to test for $\mathcal{P}$, which defines the example oracle, and the distinguishers are derived from the tester. For the tester simulation lemma, the tester plays the role of the target function, and the distinguisher family comprises low-complexity approximations to the functions to be tested for $\mathcal{P}$. In Section 3.3, we combine these lemmas to prove Theorem 1.2, and we also prove Theorem 1.3.
@@ 3.1 Simulating the Oracle
In this section, we show that the probability that $T$ accepts or rejects any particular example oracle, which is defined by a distribution $\mathcal{D}\in\Delta(\mathcal{X})$ and a target function $f:\mathcal{X}\to[0,1]$, remains roughly the same upon replacing $f$ with any $(\mathcal{R}(T),\delta)$-regular simulator $\tilde{f}$, where $\mathcal{R}(T)$ is the simple distinguisher family described in Definition 3.1 below. Specifically, $\mathcal{R}(T)$ is a family of one-way restrictions of $T$, obtained by hard-wiring fixed values for all but one of its inputs. Later, when we take $T$ to be a size-$s$ circuit, it will be clear that every distinguisher of this form is also a size-$s$ circuit.
The main result of this section, which we state formally in Lemma 3.2, is significant because of the way it facilitates the construction of additional property testers. To illustrate this informally, suppose that a property is testable by a simple circuit $T$ and that we would like to argue that some other function $\widetilde{T}$ is also a valid tester. One way to do this would be to argue that $\widetilde{T}$ behaves similarly to $T$ for all possible example oracles, but this may be difficult. Lemma 3.2 shows it suffices to check that $\widetilde{T}$ behaves similarly to $T$ for example oracles defined by low-complexity target functions $\tilde{f}$. We note, however, that the result of this section will be stated without the language of property testing, as it depends only on the function $T$.
Before proceeding with the statement and proof, we first define the relevant distinguisher class. The definition is precisely the one needed to enable a certain hyrbid argument that we plan to carry out. We state the definition in a a general manner so that it is applicable to both the actual tester $T:(\mathcal{X}\times\{0,1\})^{m}\times\{0,1\}^{\ell}\to\{0,1\}$ and its mean function $\bar{T}:(\mathcal{X}\times\{0,1\})^{m}\to[0,1]$, which averages over the choice of the random seed $r\in\{0,1\}^{\ell}$.
@@
In the special case that $T$ is Boolean-valued, each function $T_{x_{\neq i},y,r}$ has codomain $\{0,1\}$. With this definition in hand, we are now ready to state the main result of this section:
@@
The proof of Lemma 3.2 involves two simple components. The first component is a hybrid argument, similar to one used in recent work on multicalibration-based characterizations of the indistinguishability of product distributions [MPV25, DT25]. The other component is a standard transformation between distinguishers that receive a labeled or unlabeled input.
@@ 3.2 Simulating the Tester
In this section, we show that replacing $T$ with a suitable supersimulator $\widetilde{T}$ only slightly affects our chance of accepting or rejecting any sufficiently low-complexity example oracle. This promise problem, in which we assume that the example oracle has low complexity, is a natural variant of the property testing framework defined in Section 2. It is motivated by our main result from the preceding section, which showed that for some purposes, it suffices to restrict attention to oracles defined by low-complexity target functions. Later, in Section 3.3, we will show how to use $\widetilde{T}$ to establish the structured symmetry of a property under consideration.
@@
$\bar{T}(x,\tilde{y})=2^{m}\cdot\operatorname*{\mathbb{E}}_{z}\bigl[\bar{T}(x,z)\,\bm{1}[\tilde{y}=z]\,\big|\,x,t\bigr].$ (*)
@@
Finally, taking the expectation over $t$ and using the identity ()(*) yields
@@ 3.3 Constructing the Partition
In this section, we prove Theorems 1.2 and 1.3. First, we instantiate Lemma 3.2 (oracle simulation) with Lemma 2.4 (complexity-theoretic regularity). Next, we instantiate Lemma 3.4 (tester simulation) with Lemma 2.6 (supersimulators). Finally, we turn our attention to small circuit testers and use the fact that operations like restrictions and thresholding roughly preserve circuit size.
@@
4 Extensions
When combined, Theorems 1.2 and 1.3 give a characterization of properties testable from a constant number of samples. The characterization also holds for the class of properties testable from $m(n)$ samples, where $m(n)=O(\log^{(k)}(n))$ for all $k\in\mathbb{N}$. (Here, $\log^{(k)}$ denotes the $k$-fold iterated logarithm.) In fact, even for properties testable from as many as $O(\log n)$ samples, the more interesting direction of our equivalence (efficient testability implies structured symmetry, or Theorem 1.2) continues to hold.
In this section, we push further in this direction. In Sections 4.1 and 4.2, we study properties that are testable from larger sample sizes than $O(\log n)$. Finally, in Section 4.3, we sketch an argument showing that the results of this paper, although stated in the language of property testing of Boolean functions, can be easily rephrased in the language of testing high-entropy distributions over an arbitrary domain.
4.1 From $O(\log n)$- to $O(n)$-Sample Testability via Consistency Counting
The first extension we consider concerns the number of parts in the partition. It was previously known that $m$-sample testing implies symmetry with $2^{2^{O(m)}}$ parts [BY19]. (Similarly, in Section 3, we showed that efficient $m$-sample testing implies structured symmetry with $2^{2^{O(m)}}$ parts.) In this section, we exponentially improve the dependence on the tester’s sample size, reducing the number of structured sets to just $2^{O(m)}$. Phrased differently, while [BY19] and Theorem 1.2 only give nontrivial conclusions for properties testable with $O(\log n)$ samples, where $2^{n}$ is the domain size, the result of this section remains meaningful for properties testable with $O(n)$ samples.
In order to achieve this improvement, we allow overlaps among the sets that we had previously insisted form a partition. Of course, allowing overlaps requires a corresponding change to the very meaning of “symmetry.” To motivate our new definition, note that a Boolean function property is symmetric with respect to a partition if and only if it depends only on the function’s densities within the parts of the partition. Analogously, in this section, we will show that the property depends only on the function’s densities on a collection of (possibly overlapping) sets. In fact, we will prove an even stronger result: Any $m$-sample testable property essentially boils down to performing consistency counting on a collection of $2^{O(m)}$ special functions.
Definition 4.1 (Consistency Counting).
We say that a collection of $m$ samples $(x_{i},y_{i})$ is consistent with a function $f:\mathcal{X}\to\{0,1\}$ if $y_{i}=f(x_{i})$ for all indices $i\in[m]$. Given a family of good functions $\mathcal{F}_{+}$ and a family of bad functions $\mathcal{F}_{-}$, the $(m,\mathcal{F}_{+},\mathcal{F}_{-})$-consistency counter $T:(\mathcal{X}\times\{0,1\})^{m}\to\{0,1\}$ accepts inputs that are consistent with more good functions than bad.
Clearly, if we run a consistency counter on samples labeled by an unknown function, then the expected output of the tester depends only on the function’s densities on the sets $\{x\in\mathcal{X}\mid f^{\prime}(x)=1\}$, where $f^{\prime}$ is either one of the special “good” functions $\mathcal{F}_{+}$ or one of the special “bad” functions $\mathcal{F}_{-}$. Thus, consistency counting is inherently density-based. However, the converse is not necessarily true. Consistency counters are especially simple density-based testers, which just check whether each of the $k$ special functions could have plausibly labeled the $m$ observed samples. The following theorem, which is the main result of this section, states that every property that is testable from few samples has a consistency counter using a similar number of samples.
Theorem 4.2.
Note that there is a straightforward converse to Theorem 4.2: By definition, the $(m,\mathcal{F}_{+},\mathcal{F}_{-})$-consistency counter uses just $m$ samples, so if a property is testable with such a consistency counter, then it is testable with $m$ samples.
The proof of Theorem 4.2 is broadly similar to that of Theorem 1.2, but it makes additional use of the additive structure of the simulator from the complexity-theoretic regularity lemma. Since we do not consider computational complexity in this section, we will not require the supersimulators lemma like we did in the previous section. Instead, our proof will only require the basic complexity-theoretic regularity lemma.
@@ 4.1 From $O(\log n)$- to $O(n)$-Sample Testability via Consistency Counting
Next, fix $\gamma>0$. By Lemma 2.4 (complexity-theoretic regularity), there exists a function $\widetilde{T}\in\mathcal{S}_{k,\gamma/2}(\mathcal{C})$ that is a $(\mathcal{C},\gamma)$-regular simulator for $\bar{T}$, where $k<2/\gamma^{2}$. Then, by Lemma 3.4 (tester simulation), we have
@@
$\widetilde{T}=\biggl[\frac{\gamma}{2}\bigl(\sigma_{1}F_{1}+\cdots+\sigma_{k}F_{k}\bigr)\biggr]_{0}^{1}$
@@
$\widetilde{T}(x,y)=\biggl[\frac{\gamma}{2}\Bigl\lvert\big\{f\in\mathcal{F}_{+}\;\big|\;\forall i\in[m],\,y_{i}=f(x_{i})\bigr\}\Bigr\rvert-\frac{\gamma}{2}\Bigl\lvert\big\{f\in\mathcal{F}_{-}\;\big|\;\forall i\in[m],\,y_{i}=f(x_{i})\bigr\}\Bigr\rvert\biggr]_{0}^{1}.$
@@
4.2 Beyond $O(n)$-Sample Testability via Regularity Templates
The second extension we consider concerns regularity templates—a concept we define shortly. In Theorem 1.2, we studied properties that are efficiently testable from $O(\log n)$ samples, and in Theorem 4.2, we studied properties that are testable from $O(n)$ samples. What about properties that require larger sample sizes to test? In this section, we show that if a property is testable using $m$ samples and a circuit of size $s$, then the property is essentially testing for compatibility with a collection of regularity templates, each of complexity $O(m^{2}s)$. In particular, this result is nontrivial for all arbitrary (subexponential) sample sizes $m$ and circuit sizes $s$.
We are motivated by the classic characterization of testable graph properties, due to Alon et al. [AFNS06]. In that work, it was shown that a graph property is testable if and only if the property is regular-reducible, meaning that it essentially amounts to checking compatibility with a bounded collection of regularity templates. In the context of graph regularity, a regularity template is represented as a list of $\binom{k}{2}$ pairwise densities representing the densities between a partition of the graph into $k$ blocks. In the context of efficiently testing Boolean functions, we shall model a “regularity template” by a regular simulator for the function to be tested. We shall define the property of being “comptaible” with a regularity template via the usual notion of indistinguishability, as per Definition 2.2:
Definition 4.3.
Given a collection $\mathcal{T}\subseteq\{\{0,1\}^{n}\to[0,1]\}$ of functions, define the compatibility property $\mathcal{T}_{s,\delta}$ to be the set of functions that are $(s,\delta)$-indistinguishable from some function in $\mathcal{T}$.
The following result states that small-circuit testability (even for a moderate or large number of samples) reduces to testing compatibility with a size-bounded collection of regularity templates.
Theorem 4.4.
If the property $\mathcal{P}$ is testable with proximity $\varepsilon$ using $m$ samples and a circuit of size $s$, then there exists a set $\mathcal{T}$ of template circuits, each of size $O(m^{2}s)$, such that $\mathcal{P}\subseteq\mathcal{T}_{s,\frac{1}{13m}}\subseteq\mathcal{P}_{\varepsilon}$.
Proof.
We conclude this section with a partial converse to Theorem 4.4. Note that it is not a full converse because it sacrifices computational efficiency. In other words, the tester that is provides is not necessarily computable with a small circuit. The main idea behind the proof of the converse is a trick involving convex combinations, which was used in [AFNS06] to prove that compatibility with a fixed regularity template is testable. The rest of the proof is a series of standard applications of Chernoff/Hoeffding bounds.
Theorem 4.5.
Let $s\leq t$ with $t\log(t)\leq O(\varepsilon^{2}\delta^{2}2^{n})$. If $\mathcal{T}$ is a set of template circuits, each of size at most $t$, then the property $\mathcal{T}_{s,\delta}$ is testable with proximity $\varepsilon$ from $O(t\log(t)/\varepsilon^{2}\delta^{2})$ samples.
Proof.
By definition, $g$ has the property $\mathcal{T}_{s,\delta}$ if and only if there exists some template $h\in\mathcal{T}$ such that the above quantity is at most $\delta$ for all size-$s$ $f$ and $\sigma\in\{\pm 1\}$. For now, consider a fixed $h\in\mathcal{T}$. Note that there are at most $2^{O(s\log s)}$ circuits $f$ of size at most $s$. Therefore, by Hoeffding’s inequality, we can estimate the quantity displayed above for all size-$s$ functions $f$ and signs $\sigma$ up to error $\alpha$ with failure probability $\beta$, given $m=O((s\log(s)+\log(1/\beta))/\alpha^{2})$ labeled samples. Let us output Accept if all of these estimates are at most $\delta+\alpha$, and output Reject otherwise. Clearly, if $g\in\mathcal{T}_{s,\delta}$, then the tester accepts with probability with at least $1-\beta$.
4.3 From Function Testing to High-Entropy Distribution Testing
In this section, we sketch the third and final extension we consider, which concerns high-entropy, or dense, distribution testing. So far, we have focused on the problem of testing properties of Boolean functions from samples, but in some sense, this framing is too restrictive. Indeed, in this section, we show that our main result can be easily extended to testing properties of distributions on arbitrary domains, provided that the distribution to be tested is known to have high density (equivalently, high min-entropy).
@@ 4.3 From Function Testing to High-Entropy Distribution Testing
First, we review the relevant notions of density and entropy. Given a finite domain $\mathcal{X}$, fix a distribution $\mathcal{D}_{0}\in\Delta(\mathcal{X})$, which we call the base distribution or reference distribution. For example, when $\mathcal{X}=\{0,1\}^{n}$, we take $\mathcal{D}_{0}$ to be the uniform distribution on $\mathcal{X}$. Given another distribution $\mathcal{D}$ whose support is contained in that of $\mathcal{D}_{0}$, we say that $\mathcal{D}$ is $\mu$-dense in $\mathcal{D}_{0}$ if the ratio of $\mathcal{D}$ to $\mathcal{D}_{0}$ is at most $1/\mu$ pointwise. For example, when $\mathcal{X}=\{0,1\}^{n}$, a distribution is $\mu$-dense if it is a uniform distribution over a set of size $\mu\lvert\mathcal{X}\rvert$, or a convex combination thereof. In this case, we equivalently say that the distribution has min-entropy at least $k=n-\log(1/\mu)$.
@@
Dense Distribution Testing
With the aforementioned example in mind, we now consider a natural analogue of the property testing framework from Section 2 that applies not just to uniform $n$-bit strings with Boolean labels, but also more general dense distribution properties. For this, recall that the total variation distance between two distributions $\mathcal{D}_{0},\mathcal{D}_{1}\in\Delta(\mathcal{X})$ is their maximum absolute difference over any subset of the domain: $\max_{S\subseteq\mathcal{X}}\,\lvert\mathcal{D}_{0}(S)-\mathcal{D}_{1}(S)\rvert$.
@@
We now state the main result of this section, which directly generalizes Theorem 1.2 (efficient testing of Boolean functions implies structured symmetry) to the setting of $\mu$-dense distribution testing. The converse, of course, is straightforward. Since the proof of Theorem 4.7 is extremely similar to that of Theorem 1.2, which we carried out in detail in Section 3, we simply sketch the main alterations to the argument.
Theorem 4.7.
Fix $\mu\in(0,1/2)$. If a distribution property $\mathcal{P}$ is $\mu$-densely testable with proximity $\varepsilon$ using $m$ samples and a circuit of size $s$, then $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$ for some $2^{(1/\mu)^{O(m)}}$-part symmetric distribution property $\mathcal{Q}$ with partition complexity $(1/\mu)^{O(m)}s$.

### v2

@@ 3 Finding Structured Symmetry
In this section, we prove Theorems 1.2 and 1.3, which comprise our main equivalence between efficient testability and structured symmetry. As discussed in the introduction, this result strengthens the main theorem of [BY19] by capturing the relationship between the tester’s complexity and the structure of the property’s invariant sets. For convenience, we recall the statements of these two theorems below. The first, Theorem 1.2, is the harder of the two, showing that efficient testing implies structured symmetry.
@@
While the proof of Theorem 1.3 is quite simple, the proof of Theorem 1.2 is not, requiring two key lemmas. We call them the oracle simulation lemma and the tester simulation lemma, and we prove them in Sections 3.1 and 3.2, respectively. While both lemmas involve regular simulators, they differ in terms of which objects play the roles of the target function and distinguisher family. For the oracle simulation lemma, the target function is the one we wish to test for $\mathcal{P}$, which defines the sample-generating oracle, and the distinguishers are derived from the tester. For the tester simulation lemma, the tester plays the role of the target function, and the distinguisher family comprises low-complexity approximations to the functions to be tested for $\mathcal{P}$. In Section 3.3, we combine these lemmas to prove Theorem 1.2, and we also prove Theorem 1.3.
@@ 3.1 Simulating the Oracle
In this section, we consider the probability that an $m$-sample tester $T$ accepts or rejects inputs generated by any particular sample-generating oracle, which is defined by a distribution $\mathcal{D}\in\Delta(\mathcal{X})$ and a target function $f:\mathcal{X}\to[0,1]$. We show that this probability remains roughly the same upon replacing $f$ with any $(\mathcal{R}(T),\delta)$-regular simulator $\tilde{f}$, where $\mathcal{R}(T)$ is the simple distinguisher family described in Definition 3.1 below. Specifically, $\mathcal{R}(T)$ is a family of one-way restrictions of $T$, obtained by hard-wiring fixed values for all but one of its inputs. Later, when we take $T$ to be a size-$s$ circuit, it will be clear that every distinguisher of this form is also a size-$s$ circuit.
The main result of this section, which we state formally in Lemma 3.2, is significant because of the way it facilitates the construction of additional property testers. To illustrate this informally, suppose that a property is testable by a simple circuit $T$ and that we would like to argue that some other function $\widetilde{T}$ is also a valid tester. One way to do this would be to argue that $\widetilde{T}$ behaves similarly to $T$ for all possible sample-generating oracles, but this may be difficult. Lemma 3.2 shows it suffices to check that $\widetilde{T}$ behaves similarly to $T$ for sample-generating oracles defined by low-complexity target functions $\tilde{f}$. We note, however, that the result of this section will be stated without the language of property testing, as it depends only on the function $T$.
Before proceeding with the statement and proof, we first define the relevant distinguisher class. The definition is precisely the one needed to enable a certain hybrid argument that we plan to carry out. We state the definition in a general manner so that it is applicable to both the actual tester $T:(\mathcal{X}\times\{0,1\})^{m}\times\{0,1\}^{\ell}\to\{0,1\}$ and its mean function $\bar{T}:(\mathcal{X}\times\{0,1\})^{m}\to[0,1]$, which averages over the choice of the random seed $r\in\{0,1\}^{\ell}$.
@@
In the special case that $T$ is Boolean-valued, each function $T_{x_{\neq i},y,r}$ is also Boolean-valued. With this definition in hand, we are now ready to state the main result of this section:
@@
The proof of Lemma 3.2 involves two simple components. The first component is a hybrid argument, similar to one used in recent work on multicalibration-based characterizations of the indistinguishability of product distributions [MPV25]. The other component is a standard transformation between distinguishers that receive a labeled or unlabeled input.
@@ 3.2 Simulating the Tester
In this section, we show that replacing $T$ with a suitable supersimulator $\widetilde{T}$ only slightly affects our chance of accepting or rejecting any sufficiently low-complexity sample-generating oracle. This promise problem, in which we assume that the sample-generating oracle has low complexity, is a natural variant of the property testing framework defined in Section 2. It is motivated by our result from the preceding section, which showed that for some purposes, it suffices to restrict attention to oracles defined by low-complexity target functions. Later, in Section 3.3, we will show how to use $\widetilde{T}$ to establish the structured symmetry of a property under consideration.
@@
$\bar{T}(x,\tilde{y})=2^{m}\cdot\operatorname*{\mathbb{E}}_{z}\bigl[\bar{T}(x,z)\,\bm{1}[\tilde{y}=z]\,\big|\,x,t\bigr].$ (3.5)
@@
Finally, taking the expectation over $t$ and using equation (3.5) yields
@@ 3.3 Constructing the Partition
In this section, we prove Theorems 1.2 and 1.3. First, we instantiate Lemma 3.2 (oracle simulation) with Lemma 2.4 (complexity-theoretic regularity). Next, we instantiate Lemma 3.4 (tester simulation) with Lemma 2.6 (supersimulators). Finally, we turn our attention to small circuit testers and use the fact that operations like restrictions and thresholding barely increase circuit size.
@@
3.4 Consistency Counting
In this section, we present our modification to Theorem 1.1 that circumvents the $2^{2^{O(m)}}$ dependence on the tester’s sample size $m$, which allows us to make meaningful statements about sample sizes larger than $O(\log n)$. Recall that our goal is to show that any $m$-sample testable property essentially boils down to what we call consistency counting on a collection of $2^{O(m)}$ functions. We first recall the definition of consistency counting and the statement of our theorem.
See 1.4
See 1.5
The proof of Theorem 1.5 is broadly similar to that of Theorem 1.2, but it makes additional use of the additive structure of the simulator from the complexity-theoretic regularity lemma. Since we do not consider computational complexity in this section, we will not require the supersimulators lemma like we did in the previous section. Instead, our proof will only require the basic complexity-theoretic regularity lemma. Dissecting the proof further shows that we can always ensure $\mathcal{F}_{+}\subseteq\mathcal{P}$ and $\mathcal{F}_{-}\cap\mathcal{P}_{\varepsilon}=\varnothing$.
@@ 3.4 Consistency Counting
Next, fix $\gamma>0$. By Lemma 2.4 (complexity-theoretic regularity), there exists a function $\widetilde{T}\in\mathcal{S}_{k,\gamma/2}(\mathcal{C},1/2)$ that is a $(\mathcal{C},\gamma)$-regular simulator for $\bar{T}$, where $k<2/\gamma^{2}$ and $1/2$ refers to the constant function that always outputs $1/2$. Then, by Lemma 3.4 (tester simulation), we have
@@
$\widetilde{T}=\biggl[\frac{1}{2}+\frac{\gamma}{2}\bigl(\sigma_{1}F_{1}+\cdots+\sigma_{k}F_{k}\bigr)\biggr]_{0}^{1}$
@@
$\widetilde{T}(x,y)=\biggl[\frac{1}{2}+\frac{\gamma}{2}\Bigl\lvert\big\{f\in\mathcal{F}_{+}\;\big|\;\forall i\in[m],\,y_{i}=f(x_{i})\bigr\}\Bigr\rvert-\frac{\gamma}{2}\Bigl\lvert\big\{f\in\mathcal{F}_{-}\;\big|\;\forall i\in[m],\,y_{i}=f(x_{i})\bigr\}\Bigr\rvert\biggr]_{0}^{1}.$
@@
3.5 Dense Distribution Testing
@@
In this section, we discuss our extension of Theorem 1.2 from sample-testing Boolean functions to testing properties of dense distributions. First, we review the relevant notions of density and entropy. Given a finite domain $\mathcal{X}$, fix a distribution $\mathcal{D}_{0}\in\Delta(\mathcal{X})$, which we call the base distribution or reference distribution. For example, when $\mathcal{X}=\{0,1\}^{n}$, we take $\mathcal{D}_{0}$ to be the uniform distribution on $\mathcal{X}$. Given another distribution $\mathcal{D}$ whose support is contained in that of $\mathcal{D}_{0}$, we say that $\mathcal{D}$ is $\mu$-dense in $\mathcal{D}_{0}$ if the ratio of $\mathcal{D}$ to $\mathcal{D}_{0}$ is at most $1/\mu$ pointwise. For example, when $\mathcal{X}=\{0,1\}^{n}$, a distribution is $\mu$-dense if it is a uniform distribution over a set of size $\mu\lvert\mathcal{X}\rvert$, or a convex combination thereof. In this case, we equivalently say that the distribution has min-entropy at least $k=n-\log(1/\mu)$.
@@
With the aforementioned example in mind, we now define a natural analogue of the property testing framework from Section 2 that applies not just to uniform $n$-bit strings with Boolean labels, but also more general dense distribution properties. For this, recall that the total variation distance between two distributions $\mathcal{D}_{0},\mathcal{D}_{1}\in\Delta(\mathcal{X})$ is their maximum absolute difference over any subset of the domain: $\max_{S\subseteq\mathcal{X}}\,\lvert\mathcal{D}_{0}(S)-\mathcal{D}_{1}(S)\rvert$.
@@
We now recall the statement of our result on $\mu$-dense distribution testing. The converse, of course, is straightforward. Since the proof of Theorem 1.6 is extremely similar to that of Theorem 1.2, which we carried out in detail in Section 3, we simply sketch the main alterations to the argument.
See 1.6